OpenStudy (anonymous):

what is the derivative of 3^n

OpenStudy (agent0smith):

\[\Large y = 3^n\] Start by taking natural log of both sides \[\Large \ln y = \ln 3^n\] \[\Large \ln y = n \ln 3\] Now take the power of e on both sides \[\Large e^{\ln y} =e^{ n \ln 3}\] \[\Large y = e^ {n \ln 3}\] now try differentiating.

OpenStudy (agent0smith):

Remember that ln3 is just a constant. If a is a constant, derivative of e^ax is a*e^ax

OpenStudy (anonymous):

Ok got the answer: \[3^{n} \times \ln 3\]

OpenStudy (anonymous):

thanx a lot

OpenStudy (agent0smith):

No prob. You can also get it by differentiating implicitly from this step \[\Large \ln y = n \ln 3\] \[\Large \frac{ 1 }{ y } \times \frac{ dy }{ dx } = \ln 3 \] \[\Large \frac{ dy }{ dx } = \ln 3 \times y = \ln 3 \times 3^n\] (since y = 3^n)

OpenStudy (anonymous):

nice, I didn't know that we can do that, thanx again